OpenStudy (anonymous):

Rewrite the following expression with positive exponents. (50xy)-7/10....I came up with (50xy)-7/10=(50xy)10/7...Is this correct?

jimthompson5910 (jim_thompson5910):

\[\Large \left(50xy\right)^{-\frac{7}{10}} = \frac{1}{\left(50xy\right)^{\frac{7}{10}}} \]

jimthompson5910 (jim_thompson5910):

Basically, to make the exponent positive, you flip the fraction of the base.

OpenStudy (anonymous):

Oh wow you are right. Thanks I forgot that part

jimthompson5910 (jim_thompson5910):

you're welcome